Conclusions for Graham’s law experiment and classes of chemical reaction experiment.


1
Expert's answer
2021-04-13T02:07:09-0400

The rate of effusion of a gaseous substance is inversely proportional to the square root of its molar mass.


"\dfrac{\text{Rate of effusion A}}{\text{ rate of effusion B}}=\sqrt{\dfrac{M_B}{M_A}}"